# Trajectory Optimization for UAV-Based Medical Delivery with Temporal Logic Constraints and Convex Feasible Set Collision Avoidance

### v1

@@ Abstract
Index Terms:
Trajectory optimization, temporal logic constraints, convex programming, unmanned aerial vehicles.
I Introduction
@@
Unmanned aerial vehicles (UAVs) are emerging as a transformative solution for rapid and reliable medical logistics, especially in emergency scenarios where timely delivery of critical supplies—such as blood, vaccines, or organs—can directly impact patient outcomes. Traditional ground transportation is often limited by traffic congestion, geographic barriers, or disaster-related disruptions. In contrast, UAVs offer the ability to bypass such constraints, enabling swift point-to-point delivery in both urban and rural settings. The potential for autonomous aerial systems to support healthcare infrastructure is particularly promising in time-critical missions such as blood transfusion logistics, where every minute counts [1]. By optimizing their navigation and delivery efficiency, UAVs can serve as a crucial component in future smart healthcare systems.
@@
Despite their promise, deploying UAVs in dense urban environments presents significant challenges. First, UAV trajectories must respect the underlying vehicle dynamics, including acceleration, velocity, and turning limitations, to ensure feasibility and controllability in real-time flight. Second, urban settings are filled with static obstacles such as buildings, posing serious safety risks if not properly modeled and avoided [2]. Third, the mission objectives are often not limited to reaching a single location but require the UAV to visit multiple destinations—each with individual priorities and strict time windows [3, 4]. Finally, trajectory planning must reconcile these spatial and temporal requirements with safety constraints, all while remaining computationally tractable for online or near-real-time deployment. Addressing these challenges calls for a planning framework that integrates high-level task specification with low-level dynamic and environmental constraints [5].
@@
Early approaches to UAV trajectory planning have primarily focused on geometric or graph-based methods, such as A* search, rapidly-exploring random tree, and their numerous variants [6, 7]. These methods are effective in generating collision-free paths in static environments but often neglect dynamic feasibility, control limits, or time-critical objectives. Moreover, they typically rely on post-processing to smooth or adapt the path, which may violate safety or dynamic constraints when applied directly to real vehicles [6].
@@
To address the limitations of purely geometric planning, optimal control and trajectory optimization methods have been introduced. These approaches formulate trajectory generation as a continuous optimization problem, incorporating system dynamics and control constraints directly into the planning process [8, 9, 10]. While powerful, they often require carefully tuned cost functions and may struggle to represent complex high-level tasks such as sequential visits or deadline-based deliveries [9].
@@
More recently, temporal logic frameworks—especially signal temporal logic (STL)—have been adopted to encode complex task specifications in a formal and interpretable manner [11, 12]. STL enables the expression of spatial-temporal goals such as “reach a region within a certain time window,” making it suitable for high-level mission planning. However, directly incorporating STL constraints into optimization often leads to non-convex or mixed-integer formulations [13], which can be computationally expensive and difficult to scale, especially in real-time applications.
@@
To mitigate these challenges, recent research has explored convex relaxations and smooth approximations of STL semantics [14, 11]. By reformulating STL conditions using auxiliary variables and convex functions, it becomes possible to embed temporal logic specifications into a continuous optimization framework without sacrificing tractability. Nevertheless, integrating such methods with realistic dynamics and obstacle avoidance in three-dimensional urban settings remains an open challenge, particularly when multiple delivery targets and time-sensitive constraints are involved.
@@
While prior works have advanced individual aspects of UAV path planning—such as geometric routing, trajectory optimization, or temporal logic-based task encoding—few have addressed the full integration of high-level temporal specifications, low-level vehicle dynamics, and obstacle-rich environments within a unified, tractable framework. Most approaches either simplify the environment to two dimensions [14, 15], assume unconstrained dynamics [13, 16], or rely on computationally expensive non-convex formulations that are unsuitable for real-time applications [17]. In contrast, our work bridges the gap between formal mission specification and practical trajectory generation by combining temporal logic, convex obstacle avoidance, and physically consistent UAV motion planning in three-dimensional urban space.
@@
Obstacle avoidance is addressed using the convex feasible set (CFS) method [18], which iteratively constructs linearized safe regions around a reference trajectory, enabling scalable and collision-free path planning in complex 3D environments.
The entire planning problem—including STL satisfaction, obstacle avoidance, and UAV dynamics—is formulated as a standard convex optimization problem. This convex structure ensures computational tractability and allows the problem to be efficiently solved using off-the-shelf convex programming techniques [19].
@@
II Problem Formulation
II-A System model
@@
$\mathbf{v}_{t+1}=\mathbf{v}_{t}+\mathbf{a}_{t}\Delta t$
@@
$|a^{x}_{t}|\leq a_{\text{max}},\quad|a^{y}_{t}|\leq a_{\text{max}},\quad|a^{z}_{t}|\leq a_{\text{max}},\quad\forall t\in\mathbb{Z}_{[1,T-1]}$
@@
II-B Task Specification and STL Encoding
The UAV is tasked with delivering medical supplies—such as blood or plasma units—to a set of geographically distributed hospitals within a fixed planning horizon. Let $\{\mathbf{h}_{1},\mathbf{h}_{2},\dots,\mathbf{h}_{K}\}$ denote the set of hospital locations in 3D space, where each $\mathbf{h}_{k}\in\mathbb{R}^{3}$ is a target delivery waypoint. Each hospital $\mathbf{h}_{k}$ is associated with a temporal delivery window $[\tau_{k}^{\text{start}},\tau_{k}^{\text{end}}]\subseteq[1,T]$, during which the UAV must arrive sufficiently close to the destination.
@@ II-B Task Specification and STL Encoding
where $\pi$ is a predicate associated with a real-valued predicate function $\mu^{\pi}$; $\varphi$ and $\psi$ are STL formulas; negation and conjunction of formulas are denoted by $\neg$ and $\land$; the until operator is denoted by $\boldsymbol{U}_{[a,b]}$ with $a,b\in[1,T]$. We denote by $(\boldsymbol{x},t)\models\varphi$ a trajectory $\boldsymbol{x}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{T}\}$ satisfies $\varphi$ at time $t$ using the Boolean semantics of STL [11], where $(\boldsymbol{x},t)\models\pi\Leftrightarrow\mu^{\pi}(\mathbf{x}_{t})\geq 0$. We can evaluate the satisfaction of $\varphi$ quantitatively using a robustness function $\rho^{\varphi}$ such that $\varphi$ is satisfied if and only if $\rho^{\varphi}\geq 0$ [12]. In particular, we define the eventually operator $\Diamond_{[a,b]}\varphi:=\mathrm{T}\boldsymbol{U}_{[a,b]}\varphi$, where $\mathrm{T}$ is the true predicate. We have $(\boldsymbol{x},t)\models\Diamond_{[a,b]}\varphi$ if $\exists t^{\prime}\in[t+a,t+b]$ such that $(\boldsymbol{x},t^{\prime})\models\varphi$.
II-B1 Eventually Operator for Delivery Constraints
@@
$\varphi_{k}(t):=\|\mathbf{x}_{t}-\mathbf{h}_{k}\|_{2}\leq\varepsilon,\ t\in\mathbb{Z}_{[\tau_{k}^{\text{start}},\tau_{k}^{\text{end}}]}$
@@
$(\boldsymbol{x},0)\models\Diamond_{[\tau_{k}^{\text{start}},\tau_{k}^{\text{end}}]}\varphi_{k},\quad\forall k\in\mathbb{Z}_{[1,K]}$
II-B2 Robustness Semantics
@@
$\rho_{k}(t)=\varepsilon-\|\mathbf{x}_{t}-\mathbf{h}_{k}\|_{2},\ t\in\mathbb{Z}_{[\tau_{k}^{\text{start}},\tau_{k}^{\text{end}}]}$
@@
$\rho_{k}^{\Diamond}=\max\{\rho_{k}(t)\;|\;t\in\mathbb{Z}_{[\tau_{k}^{\text{start}},\tau_{k}^{\text{end}}]}\},\ \forall k\in\mathbb{Z}_{[1,K]}.$
@@ II-C Obstacle Representation
$\mathcal{O}_{m}\!=\!\left\{\mathbf{p}\!=\!\begin{bmatrix}p_{x}\\ p_{y}\\ p_{z}\end{bmatrix}\!\in\!\mathbb{R}^{3}\!\;\middle|\;\begin{aligned} &x_{m}^{\min}\leq p_{x}\leq x_{m}^{\max},\\ &y_{m}^{\min}\leq p_{y}\leq y_{m}^{\max},\\ &z_{m}^{\min}\leq p_{z}\leq z_{m}^{\max}\end{aligned}\right\},\forall m\!\in\!\mathbb{Z}_{[1,M]}$
@@
$\mathbf{x}_{t}\notin\mathcal{O}_{m},\ \forall m\in\mathbb{Z}_{[1,M]},\ \forall t\in\mathbb{Z}_{[1,T]}$
II-D Objective Function
@@
$J_{\text{control}}=\sum_{t=1}^{T-1}\|\mathbf{a}_{t}\|_{2}^{2}$
@@
Although obstacle avoidance is enforced through constraints, we include an obstacle proximity penalty term that discourages the UAV from flying too close to obstacle boundaries. Following [18], a signed distance function $\phi_{m}(\mathbf{x}_{t})$ is defined for each obstacle $\mathcal{O}_{m}$, and the proximity penalty is computed using a hinge loss:
$J_{\text{obs}}=\sum_{m=1}^{M}\sum_{t=1}^{T}\max(0,d_{\text{safe}}-\phi_{m}(\mathbf{x}_{t}))$
@@
$J=w_{1}J_{\text{STL}}+w_{2}J_{\text{control}}+w_{3}J_{\text{obs}}$
@@ III Optimization Framework
III-A Convex Reformulation of STL Constraints
@@
$:=\frac{1}{2}\left(\mu_{k}(t-1)\!+\!\rho_{k}(t)\!+\!\sqrt{(\mu_{k}(t-1)\!-\!\rho_{k}(t))^{2}\!+\!\alpha^{2}}\right)$
@@
These constraints are integrated into the overall trajectory optimization problem, allowing the UAV to satisfy delivery objectives in a differentiable and optimization-friendly manner. For each hospital $k$, the auxiliary robustness variable $\mu_{k}(t)$ tracks the satisfaction level of the STL eventually constraint over time. We aim to maximize the robustness, which corresponds to the strongest satisfaction of the time-windowed goal:
@@
III-B CFS for Obstacle Avoidance
In order to address the non-convexity of obstacle avoidance constraints, we adopt the CFS methodology, as originally proposed in [18]. The key idea is to transform the non-convex constraint set into a sequence of convex approximations centered around a reference trajectory. This allows the original trajectory optimization to be solved via an iterative convex programming procedure.
@@ III-B CFS for Obstacle Avoidance
$\phi_{m}^{c}(\mathbf{x}_{t}):=\phi_{m}(\bar{\mathbf{x}}_{t})+\nabla\phi_{m}(\bar{\mathbf{x}}_{t})(\mathbf{x}_{t}-\bar{\mathbf{x}}_{t})\geq 0,$
$\forall m\in\mathbb{Z}_{[1,M]},\ \forall t\in\mathbb{Z}_{[1,T]},$
where $\nabla$ denotes the subgradient that minimizes the cost function in the feasible set. This defines a local convex feasible set $\mathcal{F}^{(i)}\subseteq\Gamma$, which is used to solve the next convex subproblem. Iteration proceeds until convergence in cost or trajectory. Reformulating the original STL constraints into a smooth recursive structure not only avoids the computational burden of non-convex optimization but also ensures compatibility with convex solvers. This allows STL satisfaction to be enforced using standard tools without requiring integer variables or case-based logic. More importantly, the recursive formulation preserves the temporal semantics of the delivery task—ensuring that the UAV consistently increases its satisfaction measure within the specified time window—while enabling gradient-based optimization methods to efficiently find feasible and robust trajectories. This smooth convex approximation serves as the foundation for integrating STL specifications into practical trajectory planning.
III-C CFS for Obstacle Avoidance
In order to address the non-convexity of obstacle avoidance constraints, we adopt the Convex Feasible Set (CFS) methodology, as originally proposed in [18]. The key idea is to transform the non-convex constraint set into a sequence of convex approximations centered around a reference trajectory. This allows the original trajectory optimization to be solved via an iterative convex programming procedure.
$\phi_{m}(\mathbf{x})\geq 0\iff\mathbf{x}\notin\mathcal{O}_{m}$
$\phi_{m}^{c}(\mathbf{x}_{t}):=\phi_{m}(\bar{\mathbf{x}}_{t})+\nabla\phi_{m}(\bar{\mathbf{x}}_{t})(\mathbf{x}_{t}-\bar{\mathbf{x}}_{t})\geq 0,$
@@ III-C CFS for Obstacle Avoidance
III-D The Convex Optimization Problem

### v2

@@ Abstract
keywords:
[…]
1 Introduction
@@
Unmanned aerial vehicles are emerging as a transformative solution for rapid and reliable medical logistics, especially in emergency scenarios where timely delivery of critical supplies—such as blood, vaccines, or organs—can directly impact patient outcomes. Traditional ground transportation is often limited by traffic congestion, geographic barriers, or disaster-related disruptions. In contrast, UAVs offer the ability to bypass such constraints, enabling swift point-to-point delivery in both urban and rural settings. The potential for autonomous aerial systems to support healthcare infrastructure is particularly promising in time-critical missions such as blood transfusion logistics, where every minute counts (Bauer et al., 2021). By optimizing their navigation and delivery efficiency, UAVs can serve as a crucial component in future smart healthcare systems.
@@
Despite their promise, deploying UAVs in dense urban environments presents significant challenges. First, UAV trajectories must respect the underlying vehicle dynamics, including acceleration, velocity, and turning limitations, to ensure feasibility and controllability in real-time flight. Second, urban settings are filled with static obstacles such as buildings, posing serious safety risks if not properly modeled and avoided (Kuang and Liu, 2024). Third, the mission objectives are often not limited to reaching a single location but require the UAV to visit multiple destinations—each with individual priorities and strict time windows (Foehn et al., 2021; Romero et al., 2022). Finally, trajectory planning must reconcile these spatial and temporal requirements with safety constraints, all while remaining computationally tractable for online or near-real-time deployment. Addressing these challenges calls for a planning framework that integrates high-level task specification with low-level dynamic and environmental constraints (Szmuk et al., 2019).
@@
Early approaches to UAV trajectory planning have primarily focused on geometric or graph-based methods, such as A* search, rapidly-exploring random tree, and their numerous variants (Funk et al., 2023; Aggarwal and Kumar, 2020). These methods are effective in generating collision-free paths in static environments but often neglect dynamic feasibility, control limits, or time-critical objectives. Moreover, they typically rely on post-processing to smooth or adapt the path, which may violate safety or dynamic constraints when applied directly to real vehicles (Funk et al., 2023).
@@
To address the limitations of purely geometric planning, optimal control and trajectory optimization methods have been introduced. These approaches formulate trajectory generation as a continuous optimization problem, incorporating system dynamics and control constraints directly into the planning process (Chen et al., 2023; Morrell et al., 2018; Betts, 2010). While powerful, they often require carefully tuned cost functions and may struggle to represent complex high-level tasks such as sequential visits or deadline-based deliveries (Morrell et al., 2018).
@@
More recently, temporal logic frameworks, especially signal temporal logic, have been adopted to encode complex task specifications in a formal and interpretable manner (Gilpin et al., 2021; Belta and Sadraddini, 2019). STL enables the expression of spatial-temporal goals such as “reach a region within a certain time window,” making it suitable for high-level mission planning. However, directly incorporating STL constraints into optimization often leads to non-convex or mixed-integer formulations (Buyukkocak et al., 2021), which can be computationally expensive and difficult to scale, especially in real-time applications.
@@
To mitigate these challenges, recent research has explored convex relaxations and smooth approximations of STL semantics (Buyukkocak and Aksaray, 2025; Gilpin et al., 2021). By reformulating STL conditions using auxiliary variables and convex functions, it becomes possible to embed temporal logic specifications into a continuous optimization framework without sacrificing tractability. Nevertheless, integrating such methods with realistic dynamics and obstacle avoidance in three-dimensional urban settings remains an open challenge, particularly when multiple delivery targets and time-sensitive constraints are involved.
@@
While prior works have advanced individual aspects of UAV path planning—such as geometric routing, trajectory optimization, or temporal logic-based task encoding—few have addressed the full integration of high-level temporal specifications, low-level vehicle dynamics, and obstacle-rich environments within a unified, tractable framework. Most approaches either simplify the environment to two dimensions (Buyukkocak and Aksaray, 2025; Gundana and Kress-Gazit, 2021), assume unconstrained dynamics (Buyukkocak et al., 2021; Chen et al., 2025), or rely on computationally expensive non-convex formulations that are unsuitable for real-time applications (Zhang and Liu, 2022). In contrast, our work bridges the gap between formal mission specification and practical trajectory generation by combining temporal logic, convex obstacle avoidance, and physically consistent UAV motion planning in three-dimensional urban space.
@@
  2. 2.
    Obstacle avoidance is addressed using the convex feasible set method, which iteratively constructs linearized safe regions around a reference trajectory, enabling scalable and collision-free path planning in complex 3D environments.
  3. 3.
    The entire planning problem—including STL satisfaction, obstacle avoidance, and UAV dynamics—is formulated as a standard convex optimization problem. This convex structure ensures computational tractability and allows the problem to be efficiently solved using off-the-shelf convex programming techniques (Malyuta et al., 2022).
@@
2 Problem Formulation
2.1 System model
@@
$\mathbf{v}_{t+1}=\mathbf{v}_{t}+\mathbf{a}_{t}\Delta t$ (1)
@@
$|a^{x}_{t}|\leq a_{\text{max}},\quad|a^{y}_{t}|\leq a_{\text{max}},\quad|a^{z}_{t}|\leq a_{\text{max}},\quad\forall t\in\mathbb{Z}_{[1,T-1]}$ (2)
@@
2.2 Task Specification and STL Encoding
The UAV is tasked with delivering medical supplies—such as blood or plasma units—to a set of geographically distributed hospitals within a fixed planning horizon. Let $\{\mathbf{h}_{1},\mathbf{h}_{2},\dots,\mathbf{h}_{K}\}$ denote the set of hospital locations in 3D space, where each $\mathbf{h}_{k}\in\mathbb{R}^{3}$ is a target delivery waypoint. Each hospital $\mathbf{h}_{k}$ is associated with a temporal delivery window $\tau_{k}^{\text{start}},\tau_{k}^{\text{end}}]\subseteq[1,T]$, during which the UAV must arrive sufficiently close to the destination.
@@ 2.2 Task Specification and STL Encoding
where $\pi$ is a predicate associated with a real-valued predicate function $\mu^{\pi}$; $\varphi$ and $\psi$ are STL formulas; negation and conjunction of formulas are denoted by $\neg$ and $\land$; the until operator is denoted by $\boldsymbol{U}_{[a,b]}$ with $a,b\in[1,T]$. We denote by $(\boldsymbol{x},t)\models\varphi$ a trajectory $\boldsymbol{x}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{T}\}$ satisfies $\varphi$ at time $t$ using the Boolean semantics of STL (Gilpin et al., 2021), where $(\boldsymbol{x},t)\models\pi\Leftrightarrow\mu^{\pi}(\mathbf{x}_{t})\geq 0$. We can evaluate the satisfaction of $\varphi$ quantitatively using a robustness function $\rho^{\varphi}$ such that $\varphi$ is satisfied if and only if $\rho^{\varphi}\geq 0$ (Belta and Sadraddini, 2019). In particular, we define the eventually operator $\Diamond_{[a,b]}\varphi:=\mathrm{T}\boldsymbol{U}_{[a,b]}\varphi$, where $\mathrm{T}$ is the true predicate. We have $(\boldsymbol{x},t)\models\Diamond_{[a,b]}\varphi$ if $\exists t^{\prime}\in[t+a,t+b]$ such that $(\boldsymbol{x},t^{\prime})\models\varphi$.
2.2.1 Eventually Operator for Delivery Constraints
@@
$\varphi_{k}(t):=\|\mathbf{x}_{t}-\mathbf{h}_{k}\|_{2}\leq\varepsilon,\ t\in\mathbb{Z}_{[\tau_{k}^{\text{start}},\tau_{k}^{\text{end}}]}$ (3)
@@
$(\boldsymbol{x},0)\models\Diamond_{[\tau_{k}^{\text{start}},\tau_{k}^{\text{end}}]}\varphi_{k},\quad\forall k\in\mathbb{Z}_{[1,K]}$ (4)
2.2.2 Robustness Semantics
@@
$\rho_{k}(t)=\varepsilon-\|\mathbf{x}_{t}-\mathbf{h}_{k}\|_{2},\ t\in\mathbb{Z}_{[\tau_{k}^{\text{start}},\tau_{k}^{\text{end}}]}$ (5)
@@
$\rho_{k}^{\Diamond}=\max\{\rho_{k}(t),t\in\mathbb{Z}_{[\tau_{k}^{\text{start}},\tau_{k}^{\text{end}}]}\},\ \forall k\in\mathbb{Z}_{[1,K]}.$ (6)
@@ 2.3 Obstacle Representation
$\mathcal{O}_{m}\!=\!\left\{\mathbf{p}\!=\!\begin{bmatrix}p_{x}\\ p_{y}\\ p_{z}\end{bmatrix}\!\in\!\mathbb{R}^{3}\!\;\middle|\;\begin{aligned} &x_{m}^{\min}\leq p_{x}\leq x_{m}^{\max},\\ &y_{m}^{\min}\leq p_{y}\leq y_{m}^{\max},\\ &z_{m}^{\min}\leq p_{z}\leq z_{m}^{\max}\end{aligned}\right\},\forall m\!\in\!\mathbb{Z}_{[1,M]}$ (7)
@@
$\mathbf{x}_{t}\notin\mathcal{O}_{m},\ \forall m\in\mathbb{Z}_{[1,M]},\ \forall t\in\mathbb{Z}_{[1,T]}$ (8)
2.4 Objective Function
@@
$J_{\text{control}}=\sum_{t=1}^{T-1}\|\mathbf{a}_{t}\|_{2}^{2}$ (10)
@@
Although obstacle avoidance is enforced through constraints, we include an obstacle proximity penalty term that discourages the UAV from flying too close to obstacle boundaries. Following (Liu et al., 2018), a signed distance function $\phi_{m}(\mathbf{x}_{t})$ is defined for each obstacle $\mathcal{O}_{m}$, and the proximity penalty is computed using a hinge loss:
$J_{\text{obs}}=\sum_{m=1}^{M}\sum_{t=1}^{T}\max(0,d_{\text{safe}}-\phi_{m}(\mathbf{x}_{t}))$ (11)
@@
$J=w_{1}J_{\text{STL}}+w_{2}J_{\text{control}}+w_{3}J_{\text{obs}}$ (12)
@@ 3 Optimization Framework
3.1 Convex Reformulation of STL Constraints
@@
$=\frac{1}{2}\left(\mu_{k}(t-1)\!+\!\rho_{k}(t)\!+\!\sqrt{(\mu_{k}(t-1)\!-\!\rho_{k}(t))^{2}\!+\!\alpha^{2}}\right)$
@@
These constraints are integrated into the overall trajectory optimization problem, allowing the UAV to satisfy delivery objectives in a differentiable and optimization-friendly manner. For each hospital $k$, the auxiliary robustness variable $\mu_{k}(t)$ tracks the satisfaction level of the STL eventually constraint over time. We aim to maximize the robustness, which corresponds to the strongest satisfaction of the time-windowed goal
@@
3.2 CFS for Obstacle Avoidance
In order to address the non-convexity of obstacle avoidance constraints, we adopt the Convex Feasible Set methodology, as originally proposed in (Liu et al., 2018). The key idea is to transform the non-convex constraint set into a sequence of convex approximations centered around a reference trajectory. This allows the original trajectory optimization to be solved via an iterative convex programming procedure.
@@ 3.2 CFS for Obstacle Avoidance
$\phi_{m}^{c}(\mathbf{x}_{t})=\phi_{m}(\bar{\mathbf{x}}_{t})+\nabla\phi_{m}(\bar{\mathbf{x}}_{t})(\mathbf{x}_{t}-\bar{\mathbf{x}}_{t})\geq 0,$
@@
3.3 The Convex Optimization Problem